Rhiannon Corporation has bonds on the market with 13.5 years to maturity, a YTM of 7.40 percent, and a current price of $1,059. The bonds make semiannual payments.

What must the coupon rate be on these bonds?
Answer #1

Number of periods = 13.5 * 2 = 27

Rate = 7.40% / 2 = 3.7%

Price = Semi annual coupon * [1 - 1 / (1 + r)n]/ r + FV / (1 + r)n

1059 = Semi annual coupon * [1 - 1 / (1 + 0.037)27]/ 0.037 + 1000 / (1 + 0.037)27

1059 = Semi annual coupon * [1 - 0.37495] / 0.037 + 374.95019

1059 = Semi annual coupon * 16.893243 + 374.95019

Semi annual coupon = 40.4925

annual coupon = 40.4925 * 2 = 80.985

Coupon rate = (80.985 / 1000) * 100

Coupon rate = 8.10%